If log 2x + log 5y = 1, then xy =
Answer by jim_thompson5910(35256)

log(2x) + log(5y) = 1

log(2x*5y) = 1

log(10xy) = 1

log(10) + log(xy) = 1

1 + log(xy) = 1

log(xy) = 1-1

log(xy) = 0

xy = 10^0

xy = 1

--------------------------------------------------------------------------------------------------------------

If you need more help, email me at jim_thompson5910@hotmail.com

You can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------